OpenStudy (anonymous):

in triangle abc, the measure of angle b is 7 more than three times the measure of A. The measure of angle c is 38 more than the measure of angle a find the measure of each angle. what is the measure of angle A?

OpenStudy (anonymous):

a + b + c= 180 b= 3a+7 and c= a+38 put the values of a and b in eqn a+b+c=180 a+( 3a+7 )+(a+38 )=180 5a+ 45=180 5a= 180-45=135 a=135/5 = 27 a= 27 ans